Find the exact value of the following expression without using a calculator.
csc^2(2theta) if theta=7pi/6

Answer by nshah11(47)

csc^2(2θ) where θ = 7pi/6
= 1/(sin^2(7pi/3)) = 1/(√3/2)^2 = 4/3